OpenStudy (anonymous):

Evaluate 62 ÷ (3 + 9). 3 4 12 21

OpenStudy (hockeychick23):

so you need to use PEMDAS. (parentheses, exponents, multiply, divide, add, subtract) So focus on what's inside the parentheses so: 3+9=12. now we divide 62 by 12 so 62/12= 31/6 or 5 1/6

OpenStudy (anonymous):

12 then right

OpenStudy (hockeychick23):

nope, is this a multiple choice quiz? because none of the answers you wrote are correct, the correct answer is 5 1/6 or 31/6 or 5.166666666

OpenStudy (anonymous):

like the 62 is not together its like 6 then the lil 2 above the six then everthing els

OpenStudy (hockeychick23):

so 6 to the power of 2? or 6^2?

OpenStudy (anonymous):

6^2

OpenStudy (hockeychick23):

oh ok so we have 6^2 ÷ (3 + 9) By the rule of PEMDAS we solve the parentheses first so we have: 6^2/12 6^2= 36 so we have 36/12 36/12=3

OpenStudy (anonymous):

where u get 36 ?

OpenStudy (hockeychick23):

6^2= 6*6

OpenStudy (anonymous):

ohhhhh thnk u

OpenStudy (hockeychick23):

sure :)